OpenStudy (anonymous):

Let z_1=2+3i and z_2=1-i. Perform the indicated operations and write the solutions in the form a + b i. z_1/z_2= ?

OpenStudy (anonymous):

a. 5/2+5/2i b. -1/2+5/2i c. 1/2+5/2i d. -5/2+5/2i

OpenStudy (phi):

To do this, multiply top and bottom by the complex conjugate of z_2 the complex conjugate of z_2 is z_2 with its imaginary part negated: 1-(-i)= 1+i so you need to be able to multiply 2 complex numbers to finish the problem. do you know how to do that?

OpenStudy (anonymous):

I don't :(

OpenStudy (phi):

do you know how to multiply (a+b)(c+d) (using FOIL)?

OpenStudy (anonymous):

Oh, yes.

OpenStudy (phi):

what do you get for (a+b)(c+d) ?

OpenStudy (anonymous):

ac+ad+bc+bd

OpenStudy (phi):

OK. use the exact same idea for (1+i)(1-i)

OpenStudy (anonymous):

2?

OpenStudy (phi):

1-i+i-i^2 but i^2= -1 , so 1 - (-1)= 2 that is the denominator of your problem now do (2+3i)(1+i)

OpenStudy (anonymous):

-1+5i

OpenStudy (phi):

that is the top, so your answer is \[ \frac{-1+5i}{2} \] this is the same as \[ \frac{-1}{2} + \frac{5}{2} i \]

OpenStudy (phi):

So, to summarize, to do division, multiply top and bottom by the complex conjugate of the bottom number.

OpenStudy (anonymous):

Thanks phi! I really appreciate your help!